Hi, I need help with this question: two-thirds of a number is -5.66. What is the number?
Thank you!

Answer by Quadratic1600(28)

hi
If 2/3 = -5.66 than 1/3 = -2.83 so 3/3 = -2.83 x 3 = -8.49
-8.49